# STHN: Deep Homography Estimation for UAV Thermal Geo-localization with Satellite Imagery

### v1

@@ Abstract
Accurate geo-localization of Unmanned Aerial Vehicles (UAVs) is crucial for a variety of outdoor applications including search and rescue operations, power line inspections, and environmental monitoring. The vulnerability of Global Navigation Satellite Systems (GNSS) signals to interference and spoofing necessitates the development of additional robust localization methods for autonomous navigation. Visual Geo-localization (VG), leveraging onboard cameras and reference satellite maps, offers a promising solution for absolute localization. Specifically, Thermal Geo-localization (TG), which relies on image-based matching between thermal imagery with satellite databases, stands out by utilizing infrared cameras for effective night-time localization. However, the efficiency and effectiveness of current TG approaches, are hindered by dense sampling on satellite maps and geometric noises in thermal query images. To overcome these challenges, in this paper, we introduce STHN, a novel UAV thermal geo-localization approach that employs a coarse-to-fine deep homography estimation method. This method attains reliable thermal geo-localization within a 512-meter radius of the UAV’s last known location even with a challenging 11% overlap between satellite and thermal images, despite the presence of indistinct textures in thermal imagery and self-similar patterns in both spectra. Our research significantly enhances UAV thermal geo-localization performance and robustness against the impacts of geometric noises under low-visibility conditions in the wild. The code will be made publicly available.
@@ I Introduction
The increasing deployment of Unmanned Aerial Vehicles (UAVs) across a diverse range of applications, including agriculture [1], search and rescue operations [2], tracking [3], power line inspections [4], and solar power plant inspections [5], underscores the growing importance of robust UAV localization for autonomous navigation to guarantee the effective execution of these tasks. In outdoor environments, the ability to achieve absolute localization [6] is crucial as reliance on relative localization methods can lead to the accumulation of errors over time, particularly during long-time missions or in scenarios lacking loop closure detection. While Global Navigation Satellite Systems (GNSS) have become the preferred solutions to provide absolute locations, their reliability can be compromised by vulnerabilities to signal interference, jamming, and spoofing. Visual geo-localization [7, 8, 9, 10] emerges as a significant alternative solution, utilizing onboard visual sensors to facilitate absolute localization and navigation. This approach aligns captured RGB imagery, taken from nadir (top-down) or oblique views, with an existing reference map (such as a satellite RGB map), enabling accurate positioning even in GNSS-denied or degraded environments. However, this approach poses significant challenges in low-visibility or night-time environments.
@@
Figure 1: STHN framework for UAV thermal geo-localization with satellite maps. Considering only $11\%$ overlap between thermal images from UAV sensors and large-scale satellite maps ($1536~{}$\mathrm{m}$\times 1536~{}$\mathrm{m}$$), the framework achieves robust UAV localization within a $512~{}$\mathrm{m}$$ search radius.
@@
In response to these challenges, recent advancements in UAV thermal geo-localization [11] have explored an image-based matching approach that leverages an onboard thermal camera to capture nadir-view images. This approach aims to match these thermal images with the most similar satellite image crops from a database to determine the UAV’s location. However, this method encounters several drawbacks. Firstly, the localization accuracy is majorly influenced by the density of satellite image samples. Minimizing the sampling interval can enhance continuity between image crops, which improves localization accuracy, but it requires more extensive sampling. This substantially increases both the computation time due to extensive matching and the memory requirements. Additionally, the approach has limited tolerance for thermal images that are not correctly north-aligned, with geometric distortions such as rotation, resizing, and perspective transformation negatively impacting the localization accuracy.
@@
Addressing these limitations, this study introduces STHN framework (see Fig. 1) that leverages deep homography estimation techniques [12, 13, 14, 15, 16, 17, 18] to directly align thermal images with satellite maps of the local region, optimizing localization in GNSS-denied scenarios. This approach adopts a two-stage coarse-to-fine strategy: 1) Coarse alignment, which matches thermal images to large-scale satellite images with the distance of up to $512~{}$\mathrm{m}$$ from the last known position with a challenging constant overlap of $11\%$; and 2) Refinement, which crops and resizes the selected region and applies a second-stage estimation for enhanced accuracy.
@@
The main contributions of this research are outlined as follows. First, we introduce a novel satellite-thermal homography estimation approach tailored for UAV thermal geo-localization during nighttime, eliminating the dense satellite map sampling requirement of previous methods [11]. Second, we show that the Thermal Generative Module (TGM), proposed in [11], can also significantly enhance homography estimation accuracy, addressing the challenge of insufficient satellite-thermal image pairs — a unique obstacle when applying deep homography estimation within the thermal imaging domain. Third, we show that our results demonstrate superior localization accuracy over previous state-of-the-art real-time homography estimation methods and better efficiency over image-based matching methods. The proposed solutions also show promising results even with indistinct self-similar textures in thermal images. Finally, our results prove the robustness of STHN against the impact of certain geometric noises including rotation, resizing, and perspective transformation noises for thermal geo-localization. To our knowledge, this is the first deep homography estimation solution for UAV thermal geo-localization, facilitating reliable nighttime localization over long-distance outdoor flights.
@@ II Related Works
This section reviews the existing literature on UAV visual and thermal geo-localization and deep homography estimation, and highlights how our work differs from and advances beyond these previous efforts.
@@
UAV Visual and Thermal Geo-localization. UAV visual geo-localization technology has been explored by multiple works based on: 1) Template matching methods [19, 20] perform dense image alignment to optimize the image similarity measurements; 2) Traditional keypoint matching methods [10, 21] extract and match the keypoints using hand-crafted detector and descriptors, and 3) Deep-learning-based matching methods [8, 7, 22, 23, 24] utilize deep neural network [25] to generate robust matching features against environmental noises. For UAV thermal localization with nadir views, [26, 27] adopt Thermal Inertial Odometry (TIO) for navigating short-distance outdoor flights. For long-distance geo-localization, [28] uses keypoint-based visible-thermal image registration, which can also be used for geo-localization, whereas [11] employs image-based matching with generative models and domain adaptation for enhanced cross-spectral geo-localization with limited training data. Despite the efficiency of keypoint-based methods, their reliance on repeatable cross-spectral local features limits their applicability. In contrast, image-based matching methods [11, 29], free from this requirement, face challenges with exhaustive searches and high memory demands, with performances that are heavily dependent on satellite database density. Our research diverges by introducing deep homography estimation for precise satellite and thermal image alignment, presenting a novel geo-localization framework that surpasses the limitations of prior methods by eliminating the necessity for repeatable local features or exhaustive searches, thereby improving accuracy and efficiency.
@@
Deep Homography Estimation. Deep homography estimation is first proposed by [13], which uses four-corner displacement as the parametrization of homography estimation and four-corner perturbed images to train the model. [14] develops content-aware deep homography estimation against the noise from the dynamic dominant foreground. [15] employs inverse compositional Lucas-Kanade algorithms for multi-modal image alignment. In [16], the authors propose LocalTrans to conduct cross-resolution homography estimation. [12] shows an iterative process to iteratively refine the homography estimation results in real-time, whereas [17] uses a focus transformer for global and local correlation to enhance estimation performance. Considering UAV localization, [30] proposes to use an unsupervised approach with photometric consistency loss for warped aerial RGB images while requiring about $65\%$ overlap between two source images. For thermal imagery, [31] employs a multi-scale conditional GAN architecture [32] to conduct thermal-visible homography estimation. The subsequent work [33] shifts to a coarse-to-fine paradigm to further improve the estimation performance. However, the above works conduct experiments under similar scales of source images, having a minimum overlap of $25\%$ and, in rare instances, exactly $25\%$. Compared to these works, our approach adopts a coarse-to-fine paradigm coupled with iterative estimation. Specifically, we perform coarse estimation across different scales of source images to cover a broad area of the satellite map, resulting in a challenging constant $11\%$ overlap. For refinement, unlike other multi-scale approaches [12, 15, 16, 33], we crop the selected satellite region and perform estimations without the need for increased image resolution therefore obtaining enhanced efficiency.
@@
Figure 2: STHN Framework Overview: For the data preparation phase, TGM produces synthetic thermal images from unpaired satellite images, augmenting the dataset. The deep homography estimation phase employs an iterative update approach, using $F_{H}$ to predict the displacement $D_{RS\rightarrow RT}$ between thermal images and satellite maps. For refinement, the framework crops and resizes the selected region $B$, utilizing $F^{\prime}_{H}$ to fine-tune the four-corner displacement prediction for enhanced accuracy.
@@ III Methodology
III-A Thermal Generative Module
Due to limited satellite-thermal image pairs, we employ the Thermal Generative Module (TGM) introduced by [11] to enhance our training dataset with synthetic thermal images derived from satellite images. In the data preparation phase, we denote $I_{OS}$ and $I_{OT}$ as the pair of satellite and thermal images from the original dataset, and $I_{GS}$ as the satellite images without paired thermal images. We train TGM with the input $I_{OS}$ and target output $I_{OT}$ (detailed in [11]). After training TGM, we generate synthetic thermal images $I_{GT}$ using TGM and $I_{GS}$, and combine $I_{OS}$ and $I_{OT}$ to build an extended satellite-thermal dataset. We denote the quantity of actual thermal images as $N_{T}$ and those generated as $N_{G}$. We restrict our sampling from the generated dataset per epoch to $N_{T}$ instances to mitigate bias towards the generated dataset, given that $N_{T}\ll N_{G}$.
@@ III-B1 Coarse Alignment Stage
We denote $W_{S}$ as the width of input square satellite images $I_{S}$ and $W_{T}$ as that of input square thermal images $I_{T}$. For pre-processing, we resize $I_{S}$ and $I_{T}$ to $I_{RS}$ and $I_{RT}$ at the side length of $W_{R}$, and the resize ratios of $I_{RS}$ is $\alpha=W_{S}/W_{R}$. The homography estimation process is
@@
where $D_{RS\rightarrow RT}\in\mathbb{R}^{2\times 4}$ is the displacement from the four corners of $I_{RS}$ to those of $I_{RT}$ [^1] In other words, $D_{RS\rightarrow RT}$ is the four-corner displacement to align resized thermal images $I_{RT}$ to resized satellite images $I_{RS}$. and $F_{H}$ is the homograph estimation model. $F_{H}$ follows iterative estimation paradigm [12], which consists of three modules: A Convolutional Neural Network (CNN) [25] feature extractor outputs the feature map $\mathcal{F}_{RS},\mathcal{F}_{RT}\in\mathbb{R}^{C\times\frac{W_{R}}{4}\times\frac{W_{R}}{4}}$ ($C$ is the channel of feature map and is set to $256$), a correlation module outputs correlation volumes [34] $\mathbf{C}$ ($\frac{W_{R}}{4}\times\frac{W_{R}}{4}\times\frac{W_{R}}{4}\times\frac{W_{R}}{4}$) and $\mathbf{C}^{\frac{1}{2}}$ ($\frac{W_{R}}{4}\times\frac{W_{R}}{4}\times\frac{W_{R}}{8}\times\frac{W_{R}}{8}$), and an iterative homography estimator provides the update of the displacement $\Delta D_{RS\rightarrow RT}$. At iteration $k$, $D_{RS\rightarrow RT}$ is updated as
@@ III-B2 Refinement Stage
We create a bounding box $B$ that bounds the corners of thermal images warped by $D^{\prime}_{S\rightarrow T}$. We denote $D_{S\rightarrow B}\in\mathbb{R}^{2\times 4}$ as the four-corner displacement from $I_{S}$ to $B$. We crop out the region of $B$ to get $I_{B}$ at the side length of $W_{B}$ and resize it to $I_{RB}$ at the side length of $W_{R}$. The resize ratio is $\delta=W_{B}/W_{R}$. The refinement process is
@@
where $F^{\prime}_{H}$ has the same structure as $F_{H}$ with iterative updates (Equation 2) but does not share weights and $D_{RB\rightarrow RT}\in\mathbb{R}^{2\times 4}$ are four-corner displacement from $I_{RB}$ to $I_{RT}$. We set $\kappa=\delta/\alpha$ and the loss function is
@@
where $D_{k,RB\rightarrow RT}$ and $D^{gt}_{k,RB\rightarrow RT}$ are predicted and ground truth displacements, and $K_{2}$ is the number of updates in the refinement. $\kappa$ maps the displacement from the scale of $I_{RB}$ to the scale of $I_{RS}$, aligning with $\mathcal{L_{\textrm{coarse}}}$. The displacement of the refinement stage on the scale of $I_{S}$ is $D^{\prime\prime}_{S\rightarrow T}=\delta D_{RB\rightarrow RT}$. Combining the two stages’ results, we get the final displacement
$D_{S\rightarrow T}=D_{S\rightarrow B}+D^{\prime\prime}_{S\rightarrow T}.$ (6)
With $D_{S\rightarrow T}$, we use Direct Linear Transformation (DLT) [35] to solve the homography transformation matrix $H$. The transformed center coordinate $(x_{c},y_{c})$ is calculated as
$\left(x_{c},y_{c},1\right)^{\top}=H\times\left(\frac{W_{S}}{2},\frac{W_{S}}{2},1\right)^{\top},$ (7)
which is used for geo-localization. The total loss function is

### v2

@@ Abstract
Accurate geo-localization of Unmanned Aerial Vehicles (UAVs) is crucial for outdoor applications including search and rescue operations, power line inspections, and environmental monitoring. The vulnerability of Global Navigation Satellite Systems (GNSS) signals to interference and spoofing necessitates the development of additional robust localization methods for autonomous navigation. Visual Geo-localization (VG), leveraging onboard cameras and reference satellite maps, offers a promising solution for absolute localization. Specifically, Thermal Geo-localization (TG), which relies on image-based matching between thermal imagery with satellite databases, stands out by utilizing infrared cameras for effective nighttime localization. However, the efficiency and effectiveness of current TG approaches, are hindered by dense sampling on satellite maps and geometric noises in thermal query images. To overcome these challenges, we introduce STHN, a novel UAV thermal geo-localization approach that employs a coarse-to-fine deep homography estimation method. This method attains reliable thermal geo-localization within a 512-meter radius of the UAV’s last known location even with a challenging 11% size ratio between thermal and satellite images, despite the presence of indistinct textures and self-similar patterns. We further show how our research significantly enhances UAV thermal geo-localization performance and robustness against geometric noises under low-visibility conditions in the wild. The code is made publicly available.
Index Terms:
Deep Learning for Visual Perception, Aerial Systems: Applications, Localization
@@ I Introduction
The increasing deployment of Unmanned Aerial Vehicles (UAVs) across a diverse range of applications, including agriculture [1], search and rescue operations [2], tracking [3], power line inspections [4], and solar power plant inspections [5], underscores the growing importance of robust UAV localization for autonomous navigation to guarantee the effective execution of these tasks. In outdoor environments, absolute localization technology [6] is crucial as reliance on relative localization methods can cause error accumulation over time, particularly during long-time missions or in scenarios lacking loop closure detection. While Global Navigation Satellite Systems (GNSS) have become the preferred solutions, their reliability can be compromised by vulnerabilities to signal interference, jamming, and spoofing. Visual geo-localization [7, 8, 9, 10] emerges as a significant alternative solution, utilizing onboard cameras to facilitate absolute localization and navigation. This approach aligns captured RGB imagery, taken from nadir (top-down) or oblique views, with an existing reference map (such as a satellite map), enabling accurate positioning in GNSS-denied environments. However, this approach poses significant challenges in low-visibility or nighttime environments.
@@
Figure 1: STHN framework for UAV thermal geo-localization with satellite maps. This framework achieves robust UAV thermal localization with a challenging size ratio of 11% between thermal and satellite images.
@@
In response to these challenges, recent advancements in UAV thermal geo-localization [11] explore an image-based matching approach with an onboard thermal camera to match nadir-view images to satellite image crops from a database. However, this method encounters several drawbacks. Firstly, the localization accuracy is majorly influenced by the density of satellite image samples in the database. Reducing the sampling interval improves continuity between image crops and localization accuracy but increases computation time and memory usage for extensive sampling. Additionally, the approach has limited tolerance for thermal images that are not correctly north-aligned, with geometric distortions negatively impacting localization accuracy.
@@
Addressing these limitations, this study introduces Satellite-Thermal Homography Network (STHN) framework (see Fig. 1) that leverages deep homography estimation techniques [12, 13, 14, 15, 16, 17, 18] to directly align thermal images with satellite maps of the local region, optimizing localization in GNSS-denied scenarios. This approach adopts a two-stage coarse-to-fine strategy: 1) Coarse alignment, which matches small thermal images to large satellite maps within a search radius of $512~{}$\mathrm{m}$$ with a challenging constant size ratio of $11\%$; and 2) Refinement, which crops and resizes the selected region and applies a second-stage estimation for enhanced accuracy.
@@
The main contributions of this research are outlined as follows. First, we introduce for the first time a novel satellite-thermal Deep Homography Estimation (DHE) method based on an efficient coarse-to-fine approach tailored for UAV nighttime Thermal Geo-localization (TG), eliminating the dense satellite map sampling requirement of [11]. Second, we introduce the Thermal Generative Module (TGM) [11] into our DHE framework, improving the alignment between thermal and satellite images with significant scale change using limited satellite-thermal paired data. Third, we validate our approach by considering extensive and comprehensive experiments in challenging scenarios where thermal images have indistinct self-similar features on the deserts and a low overlap rate ($11\%$) with satellite images. We demonstrate the superior performance of our method over state-of-the-art real-time DHE methods and better efficiency and accuracy over image-based matching methods. Our results also demonstrate that STHN can effectively tolerate and estimate certain geometric noises including rotation, resizing, and perspective transformation noises for thermal geo-localization. To our knowledge, this is the first deep homography estimation solution for UAV thermal geo-localization, facilitating reliable nighttime localization over long-distance outdoor flights.
@@ II Related Works
UAV Visual and Thermal Geo-localization. UAV visual geo-localization technology has been explored by multiple works based on: 1) Template matching methods [19, 20] perform dense image alignment to optimize the image similarity measures; 2) Traditional keypoint matching methods [10, 21] extract and match the keypoints using hand-crafted detector and descriptors; and 3) Deep-learning-based matching methods [8, 7, 22, 23, 24] utilize deep neural network [25] to generate robust matching features against environmental noises. For UAV thermal localization with nadir views, [26, 27] adopt Thermal Inertial Odometry (TIO) for navigating short-distance outdoor flights. For long-distance geo-localization, [28] uses keypoint-based visible-thermal image registration, whereas [11] employs image-based matching with generative models and domain adaptation for enhanced cross-spectral geo-localization with limited training data. Despite the efficiency of keypoint-based methods, their reliance on repeatable cross-spectral local features limits their applicability. In contrast, image-based matching methods [11, 29], free from this requirement, face challenges with exhaustive searches and high memory demands, with performances that are heavily dependent on satellite database density. Our research diverges by introducing deep homography estimation for precise satellite and thermal image alignment, presenting a novel geo-localization framework that surpasses prior limitations by eliminating the necessity for repeatable local features or exhaustive searches, improving accuracy and efficiency.
@@
Deep Homography Estimation. Deep homography estimation is first proposed by [13], which uses four-corner displacement as the parametrization of homography estimation and four-corner perturbed images to train the model. [14] develops a content-aware deep homography estimation approach against the noise from the dynamic dominant foreground. [15] employs inverse compositional Lucas-Kanade algorithms for multi-modal image alignment. In [16], the authors propose LocalTrans to conduct cross-resolution homography estimation. [12] shows an iterative process to iteratively refine the homography estimation results in real-time, whereas [17] uses a focus transformer for global and local correlation to enhance estimation performance. Considering UAV localization, [30] proposes to use an unsupervised approach with photometric consistency loss for warped aerial RGB images while requiring about $65\%$ overlap between two source images. For thermal imagery, [31] employs a multi-scale conditional GAN architecture [32] to conduct thermal-visible homography estimation. The subsequent work [33] shifts to a coarse-to-fine paradigm to further improve the estimation performance. However, the previous works commonly require a minimum overlap of $25\%$ and, in rare instances, exactly $25\%$. Compared to these works, our approach adopts a coarse-to-fine paradigm but considers coarse estimation across images with major scale change for large search regions. This results in a challenging constant $11\%$ size ratio. For refinement, our approach differs from [12, 15, 16, 33], which typically upsample aligned images. Given the small size ratio of the thermal image, a large portion of the satellite image becomes redundant and can even hinder the refinement process. Instead, we crop the selected satellite region and perform estimation without increasing image resolution to enhance efficiency.
@@
Figure 2: STHN Framework Overview: For the data preparation phase, TGM produces synthetic thermal images from unpaired satellite images, augmenting the dataset. The deep homography estimation phase employs $F_{H}$ for the Coarse Alignment Stage by predicting the displacement $D_{RS\rightarrow RT}$ between thermal images and satellite maps. For theRefinement Stage, the framework crops and resizes the selected region $B$, utilizing $F^{\prime}_{H}$ to fine-tune the four-corner displacement prediction for enhanced accuracy.
@@ III Methodology
III-A Thermal Generative Module (TGM)
We employ TGM [11] to enhance our training dataset with synthetic thermal images derived from satellite images. In the data preparation phase, we denote $I_{OS}$ and $I_{OT}$ as the pair of satellite and 8-bit thermal images from the original dataset, and $I_{GS}$ as the satellite images without paired thermal images. We train TGM with the input $I_{OS}$ and target output $I_{OT}$ following pix2pix [34] approach. After training TGM, we generate synthetic thermal images $I_{GT}$ using TGM and $I_{GS}$, and combine $I_{OS}$ and $I_{OT}$ to build an extended satellite-thermal dataset. We denote the quantity of actual thermal images as $N_{T}$ and those generated as $N_{G}$. We restrict our sampling from the generated dataset per epoch to $N_{T}$ instances to mitigate bias towards $I_{GT}$, given that $N_{T}\ll N_{G}$.
@@ III-B1 Coarse Alignment Stage
We denote $W_{S}$ as the width of input square satellite images $I_{S}$ and $W_{T}$ as that of input square 8-bit thermal images $I_{T}$. For pre-processing, we resize $I_{S}$ and $I_{T}$ to $I_{RS}$ and $I_{RT}$ at the side length of $W_{R}$, and the resize ratios of $I_{RS}$ is $\alpha=W_{S}/W_{R}$. Then, we run the model
@@
where $D_{RS\rightarrow RT}\in\mathbb{R}^{2\times 4}$ is the displacement from the four corners of $I_{RS}$ to those of $I_{RT}$ and $F_{H}$ is the homography estimation model. In other words, $D_{RS\rightarrow RT}$ aligns $I_{RT}$ into $I_{RS}$. $F_{H}$ follows an iterative estimation paradigm [12], which consists of three modules: A Convolutional Neural Network (CNN) [25] feature extractor (multiple residual blocks with multi-layer CNNs and instance normalization) outputs the feature map $\mathcal{F}_{RS},\mathcal{F}_{RT}\in\mathbb{R}^{C\times\frac{W_{R}}{4}\times\frac{W_{R}}{4}}$ ($C=256$), a correlation module outputs correlation volumes [35] $\mathbf{C}$ ($\frac{W_{R}}{4}\times\frac{W_{R}}{4}\times\frac{W_{R}}{4}\times\frac{W_{R}}{4}$) and $\mathbf{C}^{\frac{1}{2}}$ ($\frac{W_{R}}{4}\times\frac{W_{R}}{4}\times\frac{W_{R}}{8}\times\frac{W_{R}}{8}$), and an iterative homography estimator (multi-layer CNNs with group normalization) provides updates of displacement $\Delta D_{RS\rightarrow RT}$. At iteration $k$, $D_{RS\rightarrow RT}$ is updated as
@@ III-B2 Refinement Stage
We create a bounding box $B$ that bounds the corners of thermal images warped by $D^{\prime}_{S\rightarrow T}$. We set $B$ orthogonal to the image frame to ensure complete coverage of the target region, even if the coarse alignment result has rotation or perspective transformation errors. We denote $D_{S\rightarrow B}\in\mathbb{R}^{2\times 4}$ as the four-corner displacement from $I_{S}$ to $B$. We crop out the region of $B$ to get $I_{B}$ at the side length of $W_{B}$ and resize it to $I_{RB}$ at the side length of $W_{R}$. The resize ratio is $\eta=W_{B}/W_{R}$. The refinement process is
@@
where $F^{\prime}_{H}$ has the same structure as $F_{H}$ with iterative updates (see Eq. 2) but does not share weights and $D_{RB\rightarrow RT}\in\mathbb{R}^{2\times 4}$ are four-corner displacement from $I_{RB}$ to $I_{RT}$. We set $\kappa=\eta/\alpha$ and the loss function is
@@
where $D_{k,RB\rightarrow RT}$ and $D^{gt}_{k,RB\rightarrow RT}$ are predicted and ground truth displacements, and $K_{2}$ is the number of updates in the refinement. $\kappa$ maps the displacement from the scale of $I_{RB}$ to the scale of $I_{RS}$, aligning with $\mathcal{L_{\textrm{coarse}}}$. The total loss function is
@@
The displacement of the refinement stage on the scale of $I_{S}$ is $D_{B\rightarrow T}=\eta D_{RB\rightarrow RT}$. Combining the two stages’ results, we get final displacements
$D_{S\rightarrow T}=D_{S\rightarrow B}+D_{B\rightarrow T}.$ (7)
With $D_{S\rightarrow T}$, we use Direct Linear Transformation (DLT) [36] to solve the homography matrix $H\in\mathbb{R}^{3\times 3}$. The geo-localization center coordinate $(x_{c},y_{c})$ is calculated as
$\left(x_{c},y_{c},1\right)^{\top}=H\times\left(\frac{W_{S}}{2},\frac{W_{S}}{2},1\right)^{\top}.$ (8)